OpenStudy (anonymous):

EASY QUESTION! FEELING GENEROUS! 12 is 4% of what number?

OpenStudy (anonymous):

x * 4% = 12 x * 4/100 = 12 x = \[\frac{ 12*100}{ 4 }\] x= 300